# Tagged particle fluctuations for TASEP with dynamics restricted by a moving wall

### v1

@@ 1 Introduction
[…]
More generally, in [8], a non-decreasing, not necessarily linear, function $f(t)$ is considered, such that the influence of the wall on the fluctuations of a tagged particle is restricted to one macroscopic time region. As a result, the limiting distribution of the particle is determined in the form of a variational process (see Theorem 4.4 of [8]), like the variational expression of the one-point distribution of TASEP without wall. The latter was achieved for a variety of initial conditions and is a consequence of the solution of TASEP in the KPZ fixed point paper by Matetski-Quastel-Remenik [39] (see also [36, 45, 46, 3, 18, 15, 27] for other instances of such variational formulas).
@@
The fact that the limiting distribution is given as a product of distribution functions is tightly related to the fact that the fluctuation of the (rescaled) tagged particle is asymptotically the maximum of two independent random variables. In the context of TASEP, this form of distribution occurred previously in the presence of shocks in a series of papers by Ferrari and Nejjar [22, 23, 40, 21, 24, 25, 42], see also Quastel-Remenik [43] and Bufetov-Ferrari [12].
In our situation, the product form stems from the circumstance that the values of the auxiliary tagged particle process without wall in the starting formula (see Proposition 2.1) are asymptotically independent for times at macroscopic distance. However, the physical ideas on which the study of shocks relies are related to the ideas in the present paper. For instance, for shocks one typically shows that the fluctuations acquired in any mesoscopic time close to the end time are asymptotically irrelevant. In our case, the fluctuations coming from mesoscopic times close to time zero are asymptotically irrelevant.
@@ 2 Model and main results
$\tilde{X}_{T}^{i}(\tau)\coloneqq\frac{x_{\alpha T}(\alpha_{i}T-\tilde{c}_{2}^{i}\tau T^{2/3})-\tilde{\mu}^{i}(\tau,T)}{-\tilde{c}_{1}^{i}T^{1/3}},$ (2.2)
@@
$\tilde{c}_{1}^{i}\coloneqq\frac{(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{2/3}\alpha_{i}^{1/6}}{\alpha^{1/6}},\ \tilde{c}_{2}^{i}\coloneqq\frac{2(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{1/3}\alpha_{i}^{5/6}}{\alpha^{1/3}}$ (2.3)
@@
$\tilde{\mu}^{i}(\tau,T)\coloneqq\sqrt{\alpha_{i}}(\sqrt{\alpha_{i}}-2\sqrt{\alpha})T-2\tau\frac{(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{4/3}\alpha_{i}^{1/3}}{\alpha^{1/3}}T^{2/3}.$ (2.4)
@@
[…]
This weak convergence and the resulting tightness of $(\tilde{X}_{T}^{i}(\tau))$ play a key role in the proof of the main Theorem 2.5 (see [6] for definitions and properties). We equip the space of càdlàg functions on compact intervals with the Skorokhod ($J_{1}$) topology. Considering a fixed interval $[a,b]\subseteq\mathbb{R}$, we denote this space by $\mathbb{D}([a,b])$.
@@
    $f_{0}(\beta)\coloneqq\begin{cases}\xi-\sqrt{1-\beta}(\sqrt{1-\beta}-2\sqrt{\alpha}),\ &\beta\in[0,1-\alpha),\\ \xi+\alpha,\ &\beta\in[1-\alpha,1].\end{cases}$ (2.7)
@@
  • Instead of $\alpha<\alpha_{0}<\alpha_{1}$, we can assume $\alpha<\alpha_{0}$ and for some $\sigma>0$, $\alpha_{1}T-\alpha_{0}T\geq T^{2/3+\sigma}$ for all $T$ large enough.
@@
[…]
@@
The proof of Theorem 2.5 can directly be extended to the case of $n+1$ time windows of possible wall influences, for any $n\in\mathbb{N}$. In this case, one fixes $10<\alpha<\alpha_{0}<\dots<\alpha_{n}<$ 1 and modifies Assumption 2.3 as follows: the inequality in (a) needs to hold for $t\in[0,T]$ with $|t-\alpha_{i}T|>\varepsilon T$ for $i=0,\dots,n$. In part (b), one considers $n+1$ parametrizations of $f(T-t)$ and obtains sequences $(g_{T}^{i}),i\in\{0,\dots,n\}$ with the same properties as before. The limit distribution becomes
@@
[…]
@@
    Then, we have $g_{1}(\tau)=\tau^{2}$ and $g_{0}(\tau)=\begin{cases}\tau^{2},\ &\tau\in\mathbb{R}_{-},\\ \infty,\ &\tau\in\mathbb{R}_{+}.\end{cases}$ As pointed out in Remark 2.6, the limit distribution becomes
@@
[…]
@@ 3 Backwards paths and couplings
Unless mentioned otherwise, $x(t)$ denotes a TASEP with arbitrary initial condition throughout this section. For any label $N\in\mathbb{Z}$, we define a process of backwards indices $N(t\downarrow\cdot)$ starting from time $t>0$ backwards to time 00 in the following way: we set $N(t\downarrow t)=N$. For each $s\in[0,t]$ such that [^3] By $s^{+}$, we mean a time infinitesimally larger than $s$. $N(t\downarrow s^{+})=n\in\mathbb{Z}$ and there is a suppressed jump attempt of the particle with label $n$ at time $s$, we update $N(t\downarrow s)=n-1$.
@@
Since for TASEP, the probability of simultaneous jump attempts of several particles equals 00, the backwards path $\pi_{N,t}$ almost surely has steps of size $1$.
@@ 3.1 Finite time results
Independence away from the backwards path region.
@@
For a given pair $N,t$, define the event $E_{\cal C}$ such that the backwards path $\pi_{N,t}$ starting at $x_{N}(t)$ is contained in a deterministic region $\mathcal{C}\subseteq\mathbb{Z}\times[0,t]$. Denote by $\tilde{x}(t)$ the TASEP with $\tilde{x}(0)=x(0)$ that shares its jump attempts with $x(t)$ in $\mathcal{C}$, but has a deterministic density $1$ (resp. 00) to the left (resp. right) of $\cal C$. Then
@@
In fact, supposing that in the underlying graphical construction of $x(t)$, there are no simultaneous Poisson events at any time, we actually have $E_{\cal C}\subseteq\{x_{N}(t)=\tilde{x}_{N}(t)\}$. The former event has a probability of one.
@@
Assume that the event $E_{\cal C}$ occurs. We have to show that $x_{N}(t)=\tilde{x}_{N}(t)$. This is obtained similarly as Proposition 3.4 of [21]. By definition, we have $x_{n}(t)\leq\tilde{x}_{n}(t)$ for all $n,t$. It holds $\tilde{x}_{N(t\downarrow 0)}(0)=x_{N(t\downarrow 0)}(0)$, with $N(t\downarrow\cdot)$ constructed via the evolution of $x(t)$. If we still have $N(t\downarrow\tau)=N(t\downarrow 0)$ for $\tau\in[0,t]$ and $\tilde{x}_{N(t\downarrow 0)}(\tau)$ jumps to the right, then the same holds for $x_{N(t\downarrow\tau)}(\tau)$ since otherwise, the backwards index would be updated to $N(t\downarrow 0)-1$. Thus, for $\tau_{0}$ being the time when $N(t\downarrow\tau_{0})=N(t\downarrow 0)+1$ is updated to $N(t\downarrow 0)$, we have $\tilde{x}_{N(t\downarrow\tau_{0}^{-})}(\tau_{0}^{-})=x_{N(t\downarrow\tau_{0}^{-})}(\tau_{0}^{-})$. This yields
@@
For two pairs $N_{1},t_{1}$ and $N_{2},t_{2}$ and two deterministic disjoint regions ${\cal C}_{1},{\cal C}_{2}$, define the event $E=\{\pi_{N_{1},t_{1}}\in{\cal C}_{1},\pi_{N_{2},t_{2}}\in{\cal C}_{2}\}$. Then, for any $x_{1},x_{2}\in\mathbb{Z}$,
($\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})=\mathbb{P}(x_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x_{N_{2}}(t_{2})\leq x_{2})+R$ (3.4)
@@
|($\left|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\}\cap E)\right|$ (3.5)
@@
|($|\mathbb{P}(\{x^{(1)}_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x^{(2)}_{N_{2}}(t_{2})\leq x_{2}\}\cap E)-\mathbb{P}(x^{(1)}_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x^{(2)}_{N_{2}}(t_{2})\leq x_{2})|$ (3.7)
@@
where we used that by definition of the processes, $x^{(1)}_{N_{1}}(t_{1})$ and $x^{(2)}_{N_{2}}(t_{2})$ are independent random variables. Therefore,
|($|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(x^{(1)}_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x^{(2)}_{N_{2}}(t_{2})\leq x_{2})|$ (3.8)
@@
|($\left|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(x_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x_{N_{2}}(t_{2})\leq x_{2})\right|$ (3.9)
@@
Comparison of backwards paths to TASEP with step initial condition.
@@
Next, we argue similarly for $n\in\{N(T\downarrow t_{1})+1,\dots,N\}$ and start at time $\tau_{1}$, as again for each $n\in\{N(T\downarrow t_{1})+1,\dots,N\}$ there exists a label $m\in\mathbb{N}$ such that $x_{n}(\tau_{1})=x^{\textrm{step},x_{1}}_{m}(t_{1},\tau_{1})$. ∎
@@
Claim 1: If it holds $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, then $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)-1$ is valid as well. First, suppose $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)$. We know by Lemma 3.8 that for each $t_{1}\leq s<\tau$, there is some label $m_{s}\in\mathbb{N}$ such that it holds $x_{N(T\downarrow s)}(s)=x^{\textrm{step},x_{1}}_{m_{s}}(t_{1},s)$. But then, since we premised $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, this jump and the one of the particle of $x^{\textrm{step},x_{1}}(t_{1},\cdot)$ previously at the position $x_{N(T\downarrow\tau)}(\tau)$ need to happen simultaneously at time $\tau$. As there are no simultaneous Poisson events in the construction of the processes, this cannot be. Next, suppose $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)+1$. Then, at time $\tau$ there must be a suppressed jump attempt of the $N(T\downarrow\tau)$-th particle and, since we assumed $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, also a jump attempt at the position next to the left. Again, this cannot occur. Thus, we have verified Claim 1.
Claim 2: If it holds $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)+1$, then $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)+1$ is valid as well. We notice that given a jump attempt at time $\tau$, the equality $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)$ is impossible: either the jump trial is suppressed and the backwards paths moves to the right, or the $M(t_{2}\downarrow\tau^{-})$-th particle jumps at time $\tau$. Further, if it holds $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1,$ then there need to be synchronous jump attempts at the positions $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$ and $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)$. Also this cannot occur. Thus, Claim 2 is confirmed.
@@
Comparison of increments with clock coupling.
@@
Suppose that at time $t$, the clock of the particles with label $n$ rings. Only gaps between the particles with labels $n+1$ and $n$ respectively $n$ and $n-1$ can change. There are three possible cases:
@@
The inequality (3.21) is due to the fact that whenever $\tilde{x}_{n}(t)$ jumps, the same applies to $x_{n}(t)$. ∎
@@
$y^{\textrm{step},M(t\downarrow\tau)}_{M(t\downarrow\tau)+n}(\tau;\cdot)\textrm{ and }y^{\textrm{step},\tilde{M}(t\downarrow\tau)}_{\tilde{M}(t\downarrow\tau)+n}(\tau;\cdot)$ (3.30)

### v2

@@ 1 Introduction
If $f(t)\gg t$, the effect of the wall is negligible, since the rightmost particle without the wall moves with unit speed. In this case, a tagged particle in the bulk of the system exhibits asymptotic fluctuations identical to those in the system without the wall, namely fluctuations distributed according to the GUE Tracy-Widom distribution (Johansson in [34]). On the other hand, if $f(t)\ll t$, for instance if $f(t)=vt$ with $v\in(0,1)$, the influence of the wall is relevant: in the framework of [8], the particle density has a rarefaction fan which goes over into a constant density profile until reaching the position of the wall. From the KPZ theory, one expects to still see the GUE Tracy-Widom law on the rarefaction fan, while in the constant density region one anticipates the GOE Tracy-Widom distribution as obtained for TASEP with constant non-random density as initial condition (see Sasamoto [47], Borodin-Ferrari-Prähofer [9] and Ferrari-Occelli [26] for generic density). In the transition region, the particle fluctuations should be those of the Airy2→1 process (see Borodin-Ferrari-Sasamoto [10]). These predictions actually come true as proven in Proposition 4.8 of [8].
More generally, in [8], a non-decreasing, not necessarily linear, function $f(t)$ is considered, such that the influence of the wall on the fluctuations of a tagged particle is restricted to one macroscopic time region. As a result, the limiting distribution of the particle is determined in the form of a variational formula (see Theorem 4.4 of [8]), that also gives the one-point distribution of the KPZ fixed point for a generic initial condition. This was achieved by the solution of TASEP in the KPZ fixed point paper by Matetski-Quastel-Remenik [39] (see also [36, 45, 46, 3, 18, 15, 27] for other instances of such variational formulas).
@@
Heuristically, the product structure of the limit distribution can be understood as follows. For a tagged particle to lie to the right of a given position $s$ at time $t$, all particles to its right must reach positions beyond $s$. The presence of a wall influence means that, within a time interval, these particles experience significant slowdowns, which lowers the probability of reaching the desired positions. Therefore, in the scaling limit, the distribution of the tagged particle is shaped by the dynamics of the process in the regions of wall influences. For different macroscopic times, these dynamics are asymptotically independent, leading to the observed decoupling. Thus the product form of the limiting distribution is due to the fact that the fluctuation of the (rescaled) tagged particle is asymptotically the maximum of two independent random variables. In the context of TASEP, this form of distribution occurred previously in the presence of shocks in a series of papers by Ferrari and Nejjar [22, 23, 40, 21, 24, 25, 42], see also Quastel-Remenik [43] and Bufetov-Ferrari [12].
In our situation, the product form stems from the circumstance that the values of the auxiliary tagged particle process without wall in the starting formula (see Proposition 2.1) are asymptotically independent for times at macroscopic distance. This follows from the fact that the fluctuations coming from mesoscopic times close to time zero are asymptotically irrelevant. The study of shocks relies on related physical ideas. In that case, the fluctuations acquired in any mesoscopic time close to the end time are asymptotically irrelevant.
@@ 2 Model and main results
$\tilde{X}_{T}^{i}(\tau):=\frac{x_{\alpha T}(\alpha_{i}T-\tilde{c}_{2}^{i}\tau T^{2/3})-\tilde{\mu}^{i}(\tau,T)}{-\tilde{c}_{1}^{i}T^{1/3}},$ (2.2)
@@
$\tilde{c}_{1}^{i}:=\frac{(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{2/3}\alpha_{i}^{1/6}}{\alpha^{1/6}},\ \tilde{c}_{2}^{i}:=\frac{2(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{1/3}\alpha_{i}^{5/6}}{\alpha^{1/3}}$ (2.3)
@@
$\tilde{\mu}^{i}(\tau,T):=\sqrt{\alpha_{i}}(\sqrt{\alpha_{i}}-2\sqrt{\alpha})T-2\tau\frac{(\sqrt{\alpha_{i}}-\sqrt{\alpha})^{4/3}\alpha_{i}^{1/3}}{\alpha^{1/3}}T^{2/3}.$ (2.4)
@@
For $T\to\infty$, the process $(\tilde{X}_{T}^{i}(\tau))$ converges weakly in the space of càdlàg functions on compact intervals to $(\mathcal{A}_{2}(\tau)-\tau^{2})$, where $\mathcal{A}_{2}$ denotes an Airy2 process.
This weak convergence and the resulting tightness of $(\tilde{X}_{T}^{i}(\tau))$ play a key role in the proof of the main Theorem 2.5 (see [6] for definitions and properties). We equip the space of càdlàg functions on compact intervals with the Skorokhod ($J_{1}$) topology. Considering a fixed interval $a,b]\subseteq\mathbb{R}$, we denote this space by $\mathbb{D}([a,b])$.
@@
    $f_{0}(\beta):=\begin{cases}\xi-\sqrt{1-\beta}(\sqrt{1-\beta}-2\sqrt{\alpha}),\ &\beta\in[0,1-\alpha),\\ \xi+\alpha,\ &\beta\in[1-\alpha,1].\end{cases}$ (2.7)
@@
  • Instead of $\alpha<\alpha_{0}<\alpha_{1}$ fixed (independent of $T$), we can assume $\alpha<\alpha_{0}$ fixed and for some $\sigma>0$, $\alpha_{1}T-\alpha_{0}T\geq T^{2/3+\sigma}$ for all $T$ large enough.
@@
where $\mathcal{A}_{2}^{0}$ and $\mathcal{A}_{2}^{1}$ are two independent Airy2 processes.
@@
The proof of Theorem 2.5 can directly be extended to the case of $n+1$ time windows of possible wall influences, for any $n\in\mathbb{N}$. In this case, one fixes $0<\alpha<\alpha_{0}<\dots<\alpha_{n}<1$ and modifies Assumption 2.3 as follows: the inequality in (a) needs to hold for $t\in[0,T]$ with $|t-\alpha_{i}T|>\varepsilon T$ for $i=0,\dots,n$. In part (b), one considers $n+1$ parametrizations of f(Tt)f(T-t) and obtains sequences $(g_{T}^{i}),i\in\{0,\dots,n\}$ with the same properties as before. The limit distribution becomes
@@
for Airy2 processes $\mathcal{A}_{2}^{0},\dots,\mathcal{A}_{2}^{n}$ and $S\in\mathbb{R}$.
@@
    Then, we have [^3] For $t>(1-\alpha_{0})T$, the wall constraint is removed because $f(t)=\infty$. Therefore, the convergence in Assumption 2.3 (b) for $i=0$ is restricted to $\tau\in\mathbb{R}_{-}$. $g_{1}(\tau)=\tau^{2}$ and $g_{0}(\tau)=\begin{cases}\tau^{2},\ &\tau\in\mathbb{R}_{-},\\ \infty,\ &\tau\in\mathbb{R}_{+}.\end{cases}$
    As pointed out in Remark 2.6, the limit distribution becomes
@@
    by Corollary 1.3 of [35] and Theorem 1 of [44]. By $F_{2\to 1;0}$, we denote the distribution function of the Airy2→1 process in the point 0.
@@ 3 Backwards paths and couplings
Unless mentioned otherwise, $x(t)$ denotes a TASEP with arbitrary initial condition throughout this section. For any label $N\in\mathbb{Z}$, we define a process of backwards indices $N(t\downarrow\cdot)$ starting from time $t>0$ backwards to time 0 in the following way: we set $N(t\downarrow t)=N$. For each $s\in[0,t]$ such that [^4] By $s^{+}$, we mean a time infinitesimally larger than $s$. $N(t\downarrow s^{+})=n\in\mathbb{Z}$ and there is a suppressed jump attempt of the particle with label $n$ at time $s$, we update $N(t\downarrow s)=n-1$.
@@
Since for TASEP, the probability of simultaneous jump attempts of several particles equals 0, the backwards path $\pi_{N,t}$ almost surely has steps of size $1$.
@@ 3.1 Finite time results
Independence away from the backwards path region
@@
For a given pair $N,t$, define the event $E_{\cal C}$ such that the backwards path $\pi_{N,t}$ starting at $x_{N}(t)$ is contained in a deterministic region $\mathcal{C}\subseteq\mathbb{Z}\times[0,t]$. Denote by $\tilde{x}(t)$ the TASEP with $\tilde{x}(0)=x(0)$ that shares its jump attempts with $x(t)$ in $\mathcal{C}$, but has a deterministic density $1$ (resp. 0) to the left (resp. right) of $\cal C$. Then
@@
In fact, since in the underlying graphical construction of $x(t)$, there are almost surely no simultaneous Poisson events at any time, we actually have $E_{\cal C}\subseteq\{x_{N}(t)=\tilde{x}_{N}(t)\}$ almost surely.
@@
Assume that the event $E_{\cal C}$ occurs. We have to show that $x_{N}(t)=\tilde{x}_{N}(t)$. This is obtained similarly as Proposition 3.4 of [21]. By definition, we have $x_{n}(t)\leq\tilde{x}_{n}(t)$ for all $n,t$. It holds $\tilde{x}_{N(t\downarrow 0)}(0)=x_{N(t\downarrow 0)}(0)$, with $N(t\downarrow\cdot)$ constructed via the evolution of $x(t)$. If we still have $N(t\downarrow\tau)=N(t\downarrow 0)$ for $\tau\in[0,t]$ and $\tilde{x}_{N(t\downarrow 0)}(\tau)$ jumps to the right, then the same holds for $x_{N(t\downarrow\tau)}(\tau)$ since otherwise, the backwards index would be updated to $N(t\downarrow 0)-1$. Thus, for $\tau_{0}$ being the time when $N(t\downarrow\tau_{0})=N(t\downarrow 0)+1$ is updated to $N(t\downarrow 0)$, we have [^5] By $\tau_{0}^{-}$, we mean a time infinitesimally smaller than $\tau_{0}$. $\tilde{x}_{N(t\downarrow\tau_{0}^{-})}(\tau_{0}^{-})=x_{N(t\downarrow\tau_{0}^{-})}(\tau_{0}^{-})$. This yields
@@
For two pairs $N_{1},t_{1}$ and $N_{2},t_{2}$ and two deterministic disjoint regions ${\cal C}_{1},{\cal C}_{2}$, define the event $E=\{\pi_{N_{1},t_{1}}\subseteq{\cal C}_{1},\pi_{N_{2},t_{2}}\subseteq{\cal C}_{2}\}$. Then, for any $x_{1},x_{2}\in\mathbb{Z}$,
$\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})=\mathbb{P}(x_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x_{N_{2}}(t_{2})\leq x_{2})+R$ (3.4)
@@
$\left|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\}\cap E)\right|$ (3.5)
@@
$|\mathbb{P}(\{x^{(1)}_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x^{(2)}_{N_{2}}(t_{2})\leq x_{2}\}\cap E)-\mathbb{P}(x^{(1)}_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x^{(2)}_{N_{2}}(t_{2})\leq x_{2})|$ (3.7)
@@
where we used that the processes $x^{(1)}_{N_{1}}(t_{1})$ and $x^{(2)}_{N_{2}}(t_{2})$ are independent random variables since ${\cal C}_{1}\cap{\cal C}_{2}=\varnothing$. Therefore,
$|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(x^{(1)}_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x^{(2)}_{N_{2}}(t_{2})\leq x_{2})|$ (3.8)
@@
$\left|\mathbb{P}(\{x_{N_{1}}(t_{1})\leq x_{1}\}\cap\{x_{N_{2}}(t_{2})\leq x_{2}\})-\mathbb{P}(x_{N_{1}}(t_{1})\leq x_{1})\mathbb{P}(x_{N_{2}}(t_{2})\leq x_{2})\right|$ (3.9)
@@
In the last step, we applied the same arguments as for (3.8) for both summands. ∎
@@
Comparison of backwards paths to TASEP with step initial condition
@@
Next, we argue similarly for $n\in\{N(T\downarrow t_{1})+1,\dots,N\}$ and start at time $\tau_{1}$, as again for each $n\in\{N(T\downarrow t_{1})+1,\dots,N\}$ there exists a label $m\in\mathbb{N}$ such that $x_{n}(\tau_{1})=x^{\textrm{step},x_{1}}_{m}(t_{1},\tau_{1})$. Since the set $\{N(T\downarrow t_{1}),\dots,N\}$ is finite, the iteration ends after a finite number of steps. ∎
@@
Claim 1: If it holds $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, then $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)-1$ is valid as well.
First, suppose $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)$. We know by Lemma 3.8 that for each $t_{1}\leq s<\tau$, there is some label $m_{s}\in\mathbb{N}$ such that it holds $x_{N(T\downarrow s)}(s)=x^{\textrm{step},x_{1}}_{m_{s}}(t_{1},s)$. But then, since we premised $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, this jump and the one of the particle of $x^{\textrm{step},x_{1}}(t_{1},\cdot)$ previously at the position $x_{N(T\downarrow\tau)}(\tau)$ need to happen simultaneously at time $\tau$. As there are no simultaneous Poisson events in the construction of the processes, this cannot be true. Next, suppose $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)+1$. Then, at time $\tau$ there must be a suppressed jump attempt of the $N(T\downarrow\tau)$-th particle and, since we assumed $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$, also a jump attempt at the position next to the left. Again, this cannot occur. Thus, we have verified Claim 1.
Claim 2: If it holds $x_{N(T\downarrow\tau^{-})}(\tau^{-})=x_{N(T\downarrow\tau)}(\tau)+1$, then $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)+1$ is valid as well.
We notice that given a jump attempt at time $\tau$, the equality $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)$ is impossible: either the jump trial is suppressed and the backwards paths moves to the right, or the $M(t_{2}\downarrow\tau^{-})$-th particle jumps at time $\tau$.
Further, if it holds $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau^{-})}(t_{1},\tau^{-})=x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1,$ then there needs to be a jump attempt at $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)-1$ at time $\tau$. Because of the assumption in Claim 2, there is also a jump attempt at $x^{\textrm{step},x_{1}}_{M(t_{2}\downarrow\tau)}(t_{1},\tau)=x_{N(T\downarrow\tau)}(\tau)$. But several synchronous jump attempts cannot occur. Thus, Claim 2 is confirmed.
@@
Comparison of increments with clock coupling
@@
Suppose that at time $t$, the clock of the particles with label $n$ rings. Only gaps between the particles with labels $n+1$ and $n$ respectively $n$ and n1n-1 can change. There are three possible cases:
@@
The inequality (3.21) is due to the fact that by (3.20), whenever $\tilde{x}_{n}(t)$ jumps, the same applies to $x_{n}(t)$. ∎
@@
[…]